OpenStudy (anonymous):

what is the slope of the line passing through(0,3) and (4,5) a.0.5 b.2 c.-2 d.-0.5

OpenStudy (anonymous):

(5-3)/(4-0) = 2/4 = 1/2

OpenStudy (anonymous):

0.5

OpenStudy (anonymous):

1/2

OpenStudy (across):

\[m=\frac{y_2-y_1}{x_2-x_1},\]where \(m\) is the slope and you know two points \((x_1,y_1)\) and \((x_2,y_2)\).